OpenStudy (anonymous):

let\[\huge a=10^{2 \times 1376}-10^{1376}+1\] evaluate\[\huge[2\sqrt{a}]\]

OpenStudy (anonymous):

thats integer part sign

OpenStudy (anonymous):

\[10^{1376} = x\]

OpenStudy (anonymous):

@waterineyes \[?<a<x^2\]

OpenStudy (anonymous):

Oh so the answer is \(?\).. Ha ha ha ha...

OpenStudy (anonymous):

0..

OpenStudy (anonymous):

ha ha ... lol try to finding a expression in terms of \(x\) like \((x+b)^2\) the take square root

OpenStudy (anonymous):

i mean for ?

OpenStudy (anonymous):

\[a = x^2 - x + 1 \implies (x-\frac{1}{2})^2 + \frac{3}{4}\] Like this ??

OpenStudy (anonymous):

yeah ... so what u can put instead of ?

OpenStudy (anonymous):

Meaning?? Explain more and you know about my english..

OpenStudy (anonymous):

i mean \[(x-\frac{1}{2})^2<a<x^2\]

OpenStudy (anonymous):

Yeah right...

OpenStudy (anonymous):

Now take the square root I think so it becomes: \[\pm (x - \frac{1}{2}) < a < \pm x\] Somewhat correct or not ??

OpenStudy (anonymous):

\[* \sqrt{a}\]

OpenStudy (anonymous):

Or it is just like: \[2(x - \frac{1}{2}) <2 \sqrt{a} < 2x\]

OpenStudy (anonymous):

Don't forget to tell me that I am wrong as usual...

OpenStudy (anonymous):

\(x>>0\) get rid of negative sign also

OpenStudy (anonymous):

In 2(x - 1/2) you are saying ??

OpenStudy (anonymous):

This will simply becomes 2x then..

OpenStudy (anonymous):

u got it already man \[x-\frac{1}{2}<\sqrt{a}<x\]\[2x-1<2\sqrt{a}<2x\]

OpenStudy (anonymous):

\[[2\sqrt{a}]=2\times 10^{1376}-1\]

OpenStudy (anonymous):

What is basically integer function is ???

OpenStudy (anonymous):

If [x] then x can take values less than or equal to x?? Or the values that will make it as integer ???

OpenStudy (anonymous):

sorry i was out

OpenStudy (anonymous):

[x] is the greatest integer smaller than or equal to x [2.1]=2 [-1.3]=-2

OpenStudy (anonymous):

For [-10.1] It will be: -11 ???

OpenStudy (anonymous):

yes sorry for replying late...there is something wrong with site

OpenStudy (anonymous):

Same here.. It is not working properly benim yoldashim..

OpenStudy (jiteshmeghwal9):